For a badminton doubles game, 2 players are chosen from among 5 male players and 3 female players to represent a club. In how many ways can this doubles pair be selected if 13.3.6 The team is a mixed double, comprising one male player and one female player? 13.3.7 The team is either a male pair or a female pair and no mixed pair is allowed?


1
Expert's answer
2021-05-31T18:30:16-0400

13.3.6


"\dbinom{5}{1}\dbinom{3}{1}=5(3)=15"

13.3.7


"\dbinom{5}{2}+\dbinom{3}{2}=10+3=13"